OpenStudy (ciarabelen):

What is the sum of -3x^2-7x+9 and -5x^2+6x-4?

OpenStudy (ciarabelen):

hairynmad can you help me please?

OpenStudy (anonymous):

You combine like terms. You have three different terms: x^2, x, and integer

OpenStudy (ciarabelen):

so would the x^2 + x^2 =x^4 or would it stay x^2

OpenStudy (anonymous):

No, examples 2x + 3x + 5x = 10x 2x^2 + 5x^2 = 7x^2

OpenStudy (ciarabelen):

oh ok so -8x^2-x+5?

OpenStudy (anonymous):

Your question asks for the sum, so you add both equations together. So... -3x^2-7x+9 -5x^2+6x-4 = ?

OpenStudy (anonymous):

Yes, that's the correct answer.

OpenStudy (ciarabelen):

ok thanks